### v2

# Monochromatic nonuniform hyperbolicity

Abstract.
We construct examples of continuous $\mathrm{GL}(2,\mathbb{R})$-cocycles which are not uniformly hyperbolic despite having the same non-zero Lyapunov exponents with respect to all invariant measures. The base dynamics can be any non-trivial subshift of finite type. According to a theorem of DeWitt–Gogolev and Guysinsky, such cocycles cannot be Hölder-continuous. Our construction uses the nonuniformly hyperbolic cocycles discovered by Walters in 1984.
2020 Mathematics Subject Classification:
37D25
1. Introduction
1.1. Regularity thresholds
Qualitative and quantitative properties of dynamical systems depend crucially on regularity, and rough systems may display peculiar behavior. For example, as discovered by Denjoy almost a hundred years ago, there exist non-minimal circle $C^{1}$-diffeomorphisms without periodic points, but such examples cannot be made of class $C^{2}$.
On the other hand, several perturbative techniques, such as the Closing Lemma, are known only in low regularity. As a result, the goal of describing properties of “typical” or “generic” dynamical systems becomes harder if restricted to high regularity.
Hölder continuity is a natural and frequent regularity assumption in hyperbolic dynamics: see [HK, §6.5.e] for discussion. Exotic phenomena can be found below the Hölder threshold: see [Bow, RY, Pu, Qu, AB, JM, BCS, Koc, Kos] for examples.
This paper is motivated by a result of DeWitt and Gogolev [DG] and Guysinsky [Gu]. Their theorem concerns Hölder-continuous linear cocycles over hyperbolic base dynamics and provides a strong and uniform conclusion (existence of a dominated splitting) based on a relatively weak and nonuniform hypothesis (narrow Lyapunov spectrum). In this paper, we construct examples showing that the Hölder regularity assumption cannot be dropped in the DeWitt-Gogolev-Guysinsky result.
1.2. Setting and terminology
Suppose that $X$ is a compact metric space and $T\colon X\to X$ is a homeomorphism. If $G$ is a topological group and $A\colon X\to G$ is a continuous map, then the pair $(T,A)$ is called a continuous $G$-cocycle, or simply, a $G$-cocycle. Associated to the pair $(T,A)$ we have a unique map $(x,n)\in X\times\mathbb{Z}\mapsto A^{(n)}(x)\in G$ that satisfies the following conditions:
(1.1) $A^{(1)}(x)=A(x)\,,\quad A^{(m+n)}(x)=A^{(m)}(T^{n}x)A^{(n)}(x)\,.$
Some authors call this map a cocycle (see [HK, §1.3.k]). When $G$ is a matrix group, we speak of linear cocycles.
Let us recall some basic facts about linear cocycles. We restrict our discussion to the case where $G$ is the group $\mathrm{SL}(2,\mathbb{R})$ of real $2\times 2$ matrices with determinant ±$\pm 1$. Proofs of all facts stated below can be found for example in [DF, § 3.8].
Let $(T,A)$ be $\mathrm{SL}(2,\mathbb{R})$-cocycle. Let $\mu$ be a Borel probability measure on $X$ which is invariant and ergodic with respect to $T$. Then there exists a nonnegative number $\lambda(T,A,\mu)$, called the (top) Lyapunov exponent of the cocycle $(T,A)$ with respect to the measure $\mu$, such that
(1.2) $\lim_{n\to\pm\infty}\frac{1}{n}\log\|A^{(n)}(x)\|=\lambda(T,A,\mu)\quad\text{for $\mu$-almost every $x$}\,.$
The Lyapunov spectrum of the cocycle is defined as the set
(1.3) $\Lambda(T,A)\coloneqq\big{\{}\lambda(T,A,\mu)\;\mathord{:}\;\mu\text{ is ergodic}\big{\}}\,.$
An ergodic measure is called hyperbolic if $\lambda(T,A,\mu)>0$. In that case, for $\mu$-almost every $x$, there exists a unique Oseledets splitting of $\mathbb{R}^{2}$ as a sum $E^{\mathrm{u}}(x)\oplus E^{\mathrm{s}}(x)$ of one-dimensional spaces such that $\frac{1}{n}\log\|A^{(n)}(x)v\|$ converges as $n\to\pm\infty$ to $\lambda(T,A,\mu)$ for all nonzero vectors $v$ in $E^{\mathrm{u}}(x)$, and to $-\lambda(T,A,\mu)$ for all nonzero vectors in $E^{\mathrm{s}}(x)$. The number $-\lambda(T,A,\mu)$ is the bottom Lyapunov exponent. The Oseledets splitting depends measurably on the point, and is invariant in the sense that
(1.4) $A(x)(E^{\mathrm{u}}(x))=E^{\mathrm{u}}(Tx)\,,\quad A(x)(E^{\mathrm{s}}(x))=E^{\mathrm{s}}(Tx)\,.$
We say that the cocycle $(T,A)$ is uniformly hyperbolic if the norms $\|A^{(n)}(x)\|$ grow exponentially in a uniform way, that is,
(1.5) $\liminf_{n\to\pm\infty}\inf_{x\in X}\frac{1}{|n|}\log\|A^{(n)}(x)\|>0\,.$
In that case, for every point $x\in X$, there exists a hyperbolic splitting $\mathbb{R}^{2}=E^{\mathrm{u}}(x)\oplus E^{\mathrm{s}}(x)$ such that vectors in $E^{\mathrm{s}}(x)$ (resp. $E^{\mathrm{u}}(x)$) are uniformly contracted under positive (resp. negative) iterates, that is,
(1.6) $\limsup_{n\to+\infty}\sup_{x\in X}\frac{1}{n}\log\max\left\{\big{\|}A^{(n)}(x)|_{E^{\mathrm{s}}(x)}\big{\|},\ \big{\|}A^{(-n)}(x)|_{E^{\mathrm{u}}(x)}\big{\|}\right\}<0\,.$
The hyperbolic splitting is continuous and invariant. Furthermore, it coincides $\mu$-almost everywhere with the Oseledets splitting, for any ergodic probability measure $\mu$.
We remark that all concepts introduced so far make sense in arbitrary dimension, but we will refrain from providing details. In general, there are several Lyapunov exponents, and the measure is called hyperbolic if none of them is zero. Such measures play an important role in differentiable dynamics: see [BP]. Uniform hyperbolicity was introduced in differentiable dynamics by Anosov and Smale in the 1960s. A uniformly hyperbolic splitting is always dominated in the sense of [BDV, §B.1], and the converse is true for $\mathrm{SL}(2,\mathbb{R})$-cocycles. In ODE theory, uniform hyperbolicity and dominated splittings are sometimes called exponential dichotomy [J$+$, §1.4] and exponential separation [CK, p. 189], respectively. Some types of Lyapunov spectra are studied in the papers [Boc, Pa, BS].
1.3. The main result
Back to $\mathrm{SL}(2,\mathbb{R})$-cocycles, note that if $(T,A)$ is uniformly hyperbolic, then all its ergodic measures are hyperbolic; actually, the Lyapunov spectrum (1.3) is away from zero, i.e., $\inf\Lambda(T,A)>0$. Consider the converse statement:
(1.7) $\inf\Lambda(T,A)>0\quad\overset{?}{\Rightarrow}\quad(T,A)\text{ uniformly hyperbolic}\,;$
in other words, does “uniform nonuniform hyperbolicity” imply uniform hyperbolicity? The answer is negative since there exist $\mathrm{SL}(2,\mathbb{R})$-cocycles over uniquely ergodic base dynamics which are not uniformly hyperbolic despite having a positive Lyapunov exponent with respect to the unique invariant probability measure. Herman [He] provides several examples of such cocycles. Another class of examples was constructed by Walters [Wa], and we will discuss it in detail later. Actually, the very first examples of (continuous-time) nonuniform hyperbolicity with uniquely ergodic base were exhibited in the Soviet literature: see [J$+$, §8.7] for references and a modern exposition.
Let us specialize further to the situation where $T\colon X\to X$ is a transitive subshift of finite type (SFT), and therefore (as long as we exclude the uninteresting case of finite state space), far from being uniquely ergodic. Then statement (1.7) holds for locally constant cocycle maps $A$. More generally, it holds if the cocycle map $A$ is Hölder continuous and fiber-bunched: see [Vel, Theorem 1.5]. However, statement (1.7) does not hold for all Hölder cocycles, as shown by an example of Velozo Ruiz [Vel, Theorem 4.1] (preceded by constructions of [CLR, Go]). Let us note that, even though the choice of metric in $X$ is not canonical, the class of Hölder maps is: see [KH, §1.9.a]).
The results of DeWitt and Gogolev [DG] and Guysinsky [Gu], if specialized to dimension $2$, provide other sufficient conditions for the validity of statement (1.7). Consider cocycles $(T,A)$ where $T$ is a transitive SFT and $A\colon X\to\mathrm{SL}(2,\mathbb{R})$ is a Hölder map. Assume that all ergodic measures have same Lyapunov exponent $\lambda_{0}$. That is, the Lyapunov spectrum $\Lambda(T,A)$ is a singleton $\{\lambda_{0}\}$, and thus may be called monochromatic. Also, assume that $\lambda_{0}>0$. Then, according to [DG, Theorem 1.2], the cocycle must be uniformly hyperbolic. Actually, it is sufficient to assume that the Lyapunov spectrum is close to a point: see [DG, Theorem 1.3]. (We also note that it suffices to consider the invariant measures supported on periodic orbits, due to Kalinin’s approximation theorem [Kal, Theorem 1.4], valid in the SFT / Hölder setting).
In this paper, we prove that the Hölder hypothesis in the results above is indispensable:
Theorem 1.1.
Let $T\colon X\to X$ be a subshift of finite type, where $X$ is not a finite set. Fix a constant $\lambda_{0}>0$. Then there exists a continuous map $A\colon X\to\mathrm{SL}(2,\mathbb{R})$ such that
  • $\lambda(T,A,\mu)=\lambda_{0}$ for every ergodic Borel probability measure $\mu$ and
  • the cocycle $(T,A)$ is not uniformly hyperbolic.
Actually, the conclusion holds not only for SFTs, but also for essentially any base dynamics that contains a nontrivial mixing SFT as a subsystem: see Theorem 2.9.
1.4. Localized nonuniform hyperbolicity
In our examples, the obstruction to uniform hyperbolicity is “localized” in the following sense: there exists a $T$-invariant compact set $Y\subset X$ such that $T|_{Y}$ is transitive and uniquely ergodic, the restricted cocycle $(T|_{Y},A|_{Y})$ is not uniformly hyperbolic, and for every compact invariant set $Z\subset X\smallsetminus Y$, the restricted cocycle $(T|_{Z},A|_{Z})$ is uniformly hyperbolic. Let us note that the aforementioned example of Velozo Ruiz [Vel, Theorem 4.1] has the same localization property. However, in his case the set $Y$ is relatively simple, namely the closure of an orbit which is homoclinic to a fixed point, while in our situation the set $Y$ is rather complicated: the restricted cocycle $(T|_{Y},A|_{Y})$ is of Walters type. The Katok area-preserving diffeomorphism [Kat] is another example of “localized nonuniform hyperbolicity” and in this case the set $Y$ is finite.
All those examples (including ours) have the common feature that hyperbolicity degenerates in a well-controlled way as one approaches the set $Y$. The main difficulty which the present work overcomes is allowing hyperbolicity to degenerate while keeping the Lyapunov spectrum monochromatic.
As mentioned above, our construction uses a cocycle of Walters type as an obstruction to uniform hyperbolicity. We leave the definition of Walters cocycles for the next section, but let us mention the feature that sets them apart from other examples of nonuniformly hyperbolic cocycles: Walters cocycles are formed by entrywise nonnegative matrices. This is the key property that makes our proof possible.
2. Scheme of the proof
2.1. Walters cocycles
Let us describe the class of nonuniformly hyperbolic cocycles introduced by Walters [Wa]. The base dynamics must have special properties, encapsulated in the following definition:
Definition 2.1.
Let $Y$ be a compact metric space and let $T\colon Y\to Y$ be a homeomorphism. We say that $T$ is a Veech map if $T$ is minimal and uniquely ergodic and $T^{2}$ is minimal but not uniquely ergodic.
Veech maps were first constructed by Veech in the study of a number-theoretical problem [Ve1]. His examples were later identified as being isomorphic to interval exchange transformations (IET): see [KN], [Ve2, p. 798]. We will use a different construction of Veech maps using symbolic dynamics:
Theorem 2.2.
There exists a shift space on finitely many symbols such that the corresponding shift transformation is a Veech map and has zero topological entropy.
Since we could not find an exact reference for Theorem 2.2, we will present a proof in Section 4.
The next lemma describes some basic properties of Veech maps:
Lemma 2.3.
If $T\colon Y\to Y$ is a Veech map, then $T^{2}$ admits exactly two ergodic probability measures $\nu_{0}$ and $\nu_{1}$, both of which have support $Y$. Furthermore, $T_{*}(\nu_{0})=\nu_{1}$, $T_{*}(\nu_{1})=\nu_{0}$, and the unique $T$-invariant probability measure is $\nu=\frac{\nu_{0}+\nu_{1}}{2}$.
Note that the measures $\nu_{0}$ are $\nu_{1}$ are mutually singular but “intermingled”, as they have the same support.
Proof.
Let $\nu$ be the unique $T$-invariant Borel probability measure. Since $\nu$ is $T^{2}$-invariant but $T^{2}$ is not uniquely ergodic, there exists an $T^{2}$-invariant measure $\nu_{0}$ different from $\nu$. We can assume that $\nu_{0}$ is ergodic with respect to $T^{2}$. Let $\nu_{1}\coloneqq T_{*}(\nu_{0})$. Then $T_{*}(\nu_{1})=T_{*}^{2}(\nu_{0})=\nu_{0}$. The probability measure $\frac{1}{2}\nu_{0}+\frac{1}{2}\nu_{1}$ is $T$-invariant and therefore it must coincide with $\nu$. Like $\nu_{0}$, the measure $\nu_{1}$ is ergodic with respect to $T^{2}$. Since $\nu_{0}\neq\nu$, we have $\nu_{1}\neq\nu_{0}$. So we found two different ergodic measures for $T^{2}$. Note that the ergodic decomposition of $\nu$ with respect to $T^{2}$ is $\frac{1}{2}\nu_{0}+\frac{1}{2}\nu_{1}$. Now, suppose $\tilde{\nu}_{0}$ is another ergodic measure for $T^{2}$. Then, letting $\tilde{\nu}_{1}\coloneqq T_{*}(\nu_{1})$, a repetition of the argument above yields that $\frac{1}{2}\tilde{\nu}_{0}+\frac{1}{2}\tilde{\nu}_{1}$ is an ergodic decomposition of $\nu$ with respect to $T^{2}$. By uniqueness of ergodic decomposition, we conclude that $\tilde{\nu}_{0}$ equals either $\nu_{0}$ or $\nu_{1}$. So $T^{2}$ admits exactly two ergodic measures, as we wanted to show. Since $T^{2}$ is minimal, each of those measures has full support. ∎
We follow Walters and use Veech maps to construct nonuniformly hyperbolic cocycles:
Theorem 2.4 (Walters [Wa]).
Let $T\colon Y\to Y$ be a Veech map. Then, for every $\lambda_{0}>0$, there exists a continuous map $B\colon Y\to\mathrm{SL}(2,\mathbb{R})$ such that the cocycle $(T,B)$ is not uniformly hyperbolic and its Lyapunov exponent $\lambda(T,B,\nu)$ with respect to the unique $T$-invariant probability measure $\nu$ equals $\lambda_{0}$. Furthermore, the matrices of the cocycle are of the form
(2.1) $B(y)=\begin{pmatrix}0&b(y)\\ 1/b(y)&0\end{pmatrix}\,.$
Any cocycle as in the conclusion of the theorem will be called a Walters cocycle. For the convenience of the reader, we provide a proof of Theorem 2.4. We note that the cocycle actually defined by Walters [Wa] is slightly different from ours (the determinant was not normalized to $-1$), and his proof used a different argument (se Remark 2.5 below).
Proof of Theorem 2.4.
Let $\nu_{0}$ and $\nu_{1}$ be the two $T^{2}$-invariant probability measures provided by Lemma 2.3. Since $\nu_{0}\neq\nu_{1}$, given a number $\lambda_{0}>0$, we can find a continuous function $\varphi$ such that
[…]
Let $b\coloneqq e^{\varphi}$ and define $B$ by formula (2.1). We claim that:
  • the Lyapunov exponent $\lambda(T,B,\nu)$ equals $\lambda_{0}$;
  • for $\nu_{0}$-a.e. $y$, the Oseledets spaces are $E^{\mathrm{u}}(y)=\mathbb{R}\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)$ and $E^{\mathrm{s}}(y)=\mathbb{R}\left(\begin{smallmatrix}0\\ 1\end{smallmatrix}\right)$;
  • for $\nu_{1}$-a.e. $y$, the Oseledets spaces are $E^{\mathrm{u}}(y)=\mathbb{R}\left(\begin{smallmatrix}0\\ 1\end{smallmatrix}\right)$ and $E^{\mathrm{s}}(y)=\mathbb{R}\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)$.
Consider the function
(2.3) $\psi\coloneqq\varphi-\varphi\circ T\,.$
By Lemma 2.3, $T_{*}\nu_{0}=\nu_{1}$, and so
[…]
(2.5) $B^{(2)}(y)=B(T(y))B(y)=\begin{pmatrix}e^{\psi(y)}&0\\ 0&e^{-\psi(y)}\end{pmatrix}\,;$
more generally, $B^{(2n)}(y)$ is a diagonal matrix of determinant $1$ whose top-left entry is
(2.6) $\exp\sum_{j=0}^{n-1}\psi(T^{2j}(y))\,.$
[…]
Since both measures $\nu_{0}$ and $\nu_{1}$ have full support, the Oseledets splitting cannot coincide $\nu$-a.e. with a continuous splitting. So the cocycle $(T,B)$ is not uniformly hyperbolic. ∎
Remark 2.5.
Walters [Wa] showed that the cocycle is not uniformly hyperbolic by proving that there exist points $y\in Y$ for which the sequence $\frac{1}{n}\log\|B^{(n)}(y)\|$ does not converge. In fact, the set of accumulation points of the sequence above is $[0,\lambda_{0}]$ for all points $y$ in a residual subset of $Y$; see [BPS, Corollary 4.13].
2.2. Monochromatic extensions
The key technical result of this paper is the following, whose proof is postponed to Section 3:
Theorem 2.6.
Let $T\colon X\to X$ be a biLipschitz homeomorphism of a compact metric space. Let $Y\subseteq X$ be a nonempty $T$-invariant compact set such that the restriction $T|_{Y}$ is uniquely ergodic, with a corresponding measure $\nu$. Let $B\colon X\to\mathrm{SL}(2,\mathbb{R})$ be a continuous map such that $B(x)$ is a nonnegative matrix for all $x\in X$. Assume that the Lyapunov exponent $\lambda_{0}\coloneqq\lambda(T|_{Y},B|_{Y},\nu)$ is positive. Then there exists a continuous map $A\colon X\to\mathrm{SL}(2,\mathbb{R})$ such that $A(y)=B(y)$ for all $y\in Y$ and $\lambda(T,A,\mu)=\lambda_{0}$ for every ergodic Borel probability measure $\mu$.
Remark 2.7.
We do not know if it is sufficient for the validity of the statement above to assume that the cocycle $(T|_{Y},B|_{Y})$ is monochromatic with positive exponent; our proof does use unique ergodicity. On the other hand, the biLipschitz hypothesis on $T$ is probably unnecessary, but we found it convenient for the construction.
2.3. Proof of the main results
We need one final ingredient:
Theorem 2.8 (Krieger’s embedding theorem [LM, Corollary 10.1.9]).
Let $Y$ be a shift space and let $X$ be a mixing SFT. Then $Y$ can be embedded as a proper subshift of $X$ if and only if the following two conditions are satisfied:
  1. 1.
    $h_{\mathrm{top}}(Y)<h_{\mathrm{top}}(X)$ and
  2. 2.
    for each $n\geq 1$, the number $\mathfrak{P}_{Y}(n)$ of periodic orbits with minimal period $n$ for the shift $Y$ is less than or equal to the corresponding number $\mathfrak{P}_{X}(n)$.
We now combine Theorems 2.2, 2.6, and 2.8 to prove the existence of nonuniformly hyperbolic monochromatic cocycles. We begin with the following variation of Theorem 1.1 which allows for dynamics beyond SFTs:
Theorem 2.9.
Let $T\colon X\to X$ be a biLipschitz homeomorphism of a compact metric space. Suppose that $T$ has a compact invariant set $\Lambda$ containing more than one point and such that the restriction $T|_{\Lambda}$ is topologically conjugate to a mixing subshift of finite type. Fix a constant $\lambda_{0}>0$. Then there exists a continuous map $A\colon X\to\mathrm{SL}(2,\mathbb{R})$ such that
  • $\lambda(T,A,\mu)=\lambda_{0}$ for every ergodic Borel probability measure $\mu$ and
  • the cocycle $(T,A)$ is not uniformly hyperbolic.
For example, the base dynamics $T$ can be any $C^{1}$ diffeomorphism of a compact manifold that admits a hyperbolic fixed point with a transverse homoclinic point: use [KH, Theorem 6.5.5].
Proof of Theorem 2.9.
We are given a number $\lambda_{0}>0$ and a biLipschitz homeomorphism $T\colon X\to X$ admitting an compact invariant set $\Lambda$ such that $T|_{\Lambda}$ is topologically conjugate to a mixing SFT. Furthermore, we assume that $\Lambda$ is not a singleton, which implies that $T|_{\Lambda}$ has positive topological entropy. By Theorem 2.2, there exists a shift map which is a Veech map and has zero topological entropy. Note that Veech maps have no periodic points. Therefore, by Theorem 2.8, we can find a compact invariant set $Y\subset\Lambda$ such that $T|_{Y}$ is a Veech map. Use Theorem 2.4 to construct a Walters cocycle $B\colon Y\to\mathrm{SL}(2,\mathbb{R})$ with Lyapunov exponent $\lambda_{0}$. Recall that $B(y)$ is the nonnegative antidiagonal matrix (2.1) where $b\colon Y\to\mathbb{R}$ is a continuous function. Tietze extension theorem provides a continuous extension of $b$ to the whole space $X$, which we denote by the same letter $b$. Then formula (2.1) defines a map $B\colon X\to\mathrm{SL}(2,\mathbb{R})$ that takes nonnegative values. Now we can apply Theorem 2.6 and find a continuous $A\colon X\to\mathrm{SL}(2,\mathbb{R})$ such that $A|_{Y}=B|_{Y}$ and the cocycle $(T,A)$ is monochromatic, i.e. $\lambda(T,A,\mu)=\lambda_{0}$ for every ergodic Borel probability measure $\mu$. The cocycle $(T,A)$ is not uniformly hyperbolic, since its restriction to $Y$ is a Walters cocycle. ∎
Next, we prove Theorem 1.1, stated in the Introduction.
Proof of Theorem 1.1.
Now we are given $\lambda_{0}>0$ and a SFT $T\colon X\to X$ where $X$ is an infinite set. Consider a spectral decomposition, that is, a decomposition of $X$ into disjoint nonempty compact subsets
(2.7) $X=\bigsqcup_{i=1}^{m}X_{i}\,,\quad X_{i}=\bigsqcup_{j=1}^{n_{i}}X_{i,j}\quad\text{such that}\quad T(X_{i,j})=X_{i,j+1\bmod n_{i}}$
and each $T^{n_{i}}|_{X_{i,1}}$ is mixing. Reindexing if necessary, we can assume that $X_{1,1}$ is not a singleton.
Just like in the previous proof, we can find a compact invariant set $Y\subset X_{1,1}$ such that $T^{n_{1}}|_{Y}$ is a Veech map, say with invariant measure $\nu$. Then we can find a Walters cocycle $(T^{n_{1}}|_{Y},B)$ whose Lyapunov exponent with respect to $\nu$ equals $n_{1}\lambda_{0}$. Let $\hat{B}\colon X\to\mathrm{SL}(2,\mathbb{R})$ be a continuous extension of the map $B$ such that $\hat{B}(x)$ is a nonpositive antidiagonal matrix if $x\in X_{1,1}$ and is the identity matrix otherwise. Note that the restriction of $T$ to the invariant set $\hat{Y}\coloneqq\bigcup_{j=0}^{n_{1}-1}T^{j}(Y)$ is uniquely ergodic, with corresponding measure $\hat{\nu}\coloneqq\frac{1}{n_{1}}\sum_{j=0}^{n_{1}-1}T^{j}_{*}\nu$. Furthermore, $\lambda(T,\hat{B},\hat{\nu})=\lambda_{0}$. The cocycle $(T|_{\hat{Y}},\hat{B}|_{\hat{Y}})$ is not uniformly hyperbolic, because $(T^{n_{1}}|_{Y},B)$ is not. Applying Theorem 2.6, we find a monochromatic cocycle $(T,A)$, where $A$ coincides with $\hat{B}$ on the set $\hat{Y}$. In particular, the cocycle $(T,A)$ cannot be uniformly hyperbolic. ∎
Our main results are proved, modulo Theorems 2.6 and 2.2, whose proofs are presented in the next two (entirely independent) sections.
3. Perturbing nonnegative cocycles
In this section, we will prove Theorem 2.6. Let us provide some highlights of the proof. We define an auxiliary cocycle map $P$ (see formula (3.21)) which coincides with $B$ on the set $Y$, and is formed by strictly positive matrices on its complement $X\smallsetminus Y$. We call $P$ the positive perturbation of $B$. The off-diagonal entries of the matrix $P(x)$ have a very rough modulus of regularity near $Y$. As a consequence, the cocycle $(T,P)$ admits a continuous invariant field of lines on the complement of $Y$ (see Lemma 3.8). Using unique ergodicity and positivity, we can control finite-time expansion rates near $Y$ (see Lemma 3.5). This allows us to solve an “approximate” cohomological equation for the expansion rate along the invariant bundle on the noncompact set $X\smallsetminus Y$. This equation, despite degenerating as we approach $Y$, has a vanishing error term (see Lemma 3.12). Once in possession of this information, it becomes easy to define a last perturbation $A$ (see equation (3.94)) which will have the monochromatic property.
Conventions for this section: We will use $(\epsilon_{k})$ to denote any sequence of positive numbers that converges to 00; the specific value of the sequence may change with each occurrence. Similarly, $C$ will denote any positive constant.
3.1. Fixing the time scales
For the remainder of this section, let $X$, $T$, $Y$, $\nu$, $B$, and $\lambda_{0}$ be as in the statement of Theorem 2.6.
Lemma 3.1.
We have the following limits:
(3.1) $\lambda_{0}$ $=\lim_{n\to\infty}\frac{1}{n}\int_{Y}$ $\log\|B^{(n)}(y)\|\,d\nu(y)\,,$
(3.2) $\lambda_{0}$ $=\lim_{n\to\infty}\frac{1}{n}\sup_{y\in Y}$ $\log\|B^{(n)}(y)\|\,.$
Proof.
The first limit comes follows the subadditive ergodic theorem [DF, Theorem 3.4.2], while the second one follows from the semi-uniform subadditive ergodic theorem [Mo, Theorem A.3]. ∎
The next step is to fix a superexponential sequence $(m_{k})$ of time scales along which we have some specific control of norms.
Lemma 3.2 (Time scales).
There exists a sequence of positive integers $(m_{k})_{k\geq 1}$ such that
(3.3) $\frac{m_{k+1}}{m_{k}}\nearrow\infty\,,$
and for all $y\in Y$ and all $n\geq m_{k+1}$, there exist integers $q$, $r$, $s$ such that
(3.4) $n=r+qm_{k}+s\,,\quad q>0\,,\quad|r|<m_{k}\,,\quad|s|<m_{k}\,,\quad\text{and}$
(3.5) $\frac{1}{qm_{k}}\sum_{j=0}^{q-1}\log\|B^{(m_{k})}(T^{r+jm_{k}}y)\|>\lambda_{0}-\epsilon_{k}\,.$
Proof.
The sequence $(m_{k})$ will be constructed recursively. It will have the following extra property:
(3.6) $\int_{Y}\frac{1}{m_{k}}\log\|B^{(m_{k})}\|\,d\nu>\lambda_{0}-\frac{1}{k}\,.$
Using (3.1), we can choose $m_{1}$ such that (3.6) holds with $k=1$. Now fix $k$ and as an induction hypothesis suppose that a number $m_{k}$ with property (3.6) is already chosen. Since $T|_{Y}$ is uniquely ergodic, the Birkhoff averages of the function $\frac{1}{m_{k}}\log\|B^{(m_{k})}\|$ converge uniformly to its integral with respect to the unique invariant probability $\nu$. Therefore, we can find $m_{k+1}$ such that for all $N\geq\frac{m_{k+1}}{2}$ and $y\in Y$, we have
(3.7) $\frac{1}{N}\sum_{i=0}^{N-1}\frac{1}{m_{k}}\log\|B^{(m_{k})}(T^{i}y)\|>\lambda_{0}-\frac{2}{k}\,.$
Increasing $m_{k+1}$ if necessary, we assume that $m_{k+1}>(k+1)m_{k}$ and furthermore property (3.6) holds with $k+1$ in the place of $k$ (which is feasible because of (3.1)). This completes the inductive defintion of the sequence $(m_{k})$.
Now let us check that the sequence has the desired properties. Fix $k\geq 1$, $y\in Y$, and $n\geq m_{k+1}$. Let $q\coloneqq\left\lfloor\frac{n}{m_{k}}\right\rfloor$ and $N\coloneqq qm_{k}$. Note that $N>n-m_{k}\geq\frac{m_{k+1}}{2}$, so inequality (3.7) holds. We rewrite the left hand side of this inequality as follows:
(3.8) $\frac{1}{N}\sum_{i=0}^{N-1}\frac{1}{m_{k}}\log\|B^{(m_{k})}(T^{i}y)\|=\frac{1}{m_{k}}\sum_{r=0}^{m_{k}-1}\underbrace{\frac{1}{qm_{k}}\sum_{j=0}^{q-1}\log\|B^{(m_{k})}(T^{r+jm_{k}}y)\|}_{\eqcolon a_{r}}\,.$
Choose and fix some $r\in\ldbrack 0,m_{k}-1\rdbrack$ such that $a_{r}>\lambda_{0}-\frac{2}{k}$. That is, property (3.5) holds where $\epsilon_{k}$ is actually $\frac{2}{k}$. Finally, let $s\coloneqq(n-N)-r$. Since both $n-N$ and $r$ belong to $\ldbrack 0,m_{k}-1\rdbrack$, we have $|s|<m_{k}$. Thus relations (3.4) hold. ∎
3.2. Fixing the space scales
Rescaling the metric if necessary, we assume that
(3.9) $\operatorname{diam}X<\tfrac{1}{10}\,.$
The next step is to fix a sequence $(\delta_{k})_{k\geq 1}$ of space scales associated to the sequence of time scales $(m_{k})$. Essentially, we need that $\delta_{k}\searrow 0$, and that this convergence is extremely fast. More precisely, we impose the following conditions: for each $k$,
(3.10) $\delta_{k}$ $<\tfrac{1}{10}\,,$
(3.11) $\delta_{k}$ $<e^{-e^{m_{k-4}}}\text{ if }k>4\,,$
(3.12) $\delta_{k+1}$ $<\delta_{k}e^{-k\,m_{k+1}}\,.$
Consider the sequence of Bowen metrics defined as follows:
(3.13) $\mathsf{d}_{n}(x_{1},x_{2})\coloneqq\max_{-n\leq i\leq n}\mathsf{d}(T^{i}(x_{1}),T^{i}(x_{2}))\,.$
By uniform continuity of $T^{\pm 1}$, each Bowen metric $\mathsf{d}_{n}$ is uniformly comparable to the initial metric $\mathsf{d}$. Therefore some additional care in the choice of space scales ensures the following property:
(3.14) $x_{1},x_{2}\in X,\ k>1,\ \mathsf{d}(x_{1},x_{2})\leq\delta_{k}\ \Rightarrow\ \mathsf{d}_{m_{k+1}}(x_{1},x_{2})<\delta_{k-1}\,.$
The last imposition on the sequence $(\delta_{k})$ is the following property:
[…]
which can be enforced simply because the matrix map $B$ is uniformly continuous.
3.3. Positive perturbation
In the next lemma, we fix an auxiliary continuous function $\phi$ with a very rough modulus of continuity.
Lemma 3.3.
There exists an increasing homeomorphism $\phi\colon[0,\frac{1}{10}]\to[0,\frac{1}{2}]$ such that
(3.16) $\phi(\delta_{k})$ $=e^{-m_{k-4}}$ for each $\text{for each }k>4\quad\text{and}$
(3.17) $\phi(t)$ $>\frac{1}{-\log t}$ for every $\text{for every }t\in(0,\tfrac{1}{10}]\,.$
Proof.
Let $\psi$ be the homeomorphism from $[0,\tfrac{1}{\log 10}]$ to $[0,\tfrac{1}{2}]$ defined as follows. We fix the following anchor points:
(3.18) $\psi(0)\coloneqq 0\,,\quad\psi\big{(}\tfrac{1}{\log 10}\big{)}\coloneqq\tfrac{1}{2}\,,\quad\psi\big{(}\tfrac{1}{-\log\delta_{k}}\big{)}\coloneqq e^{-m_{k-4}}\text{ for all }k>4\,;$
then we interpolate linearly between any two consecutive anchor points. It follows from (3.11) that $\psi(s)>s$ for all $s\in(0,\tfrac{1}{\log 10}]$. Then the function $\phi(t)\coloneqq\psi\big{(}\tfrac{1}{-\log t}\big{)}$ has all the desired properties. ∎
Define the following continuous functions of $x\in X$:
(3.19) $\theta(x)$ $\coloneqq\phi(\mathsf{d}(x,Y))\,,$
(3.20) $H(x)$ $\coloneqq\begin{pmatrix}\cosh\theta(x)&\sinh\theta(x)\\ \sinh\theta(x)&\cosh\theta(x)\end{pmatrix}\,,\quad\text{and}$
(3.21) $P(x)$ $\coloneqq B(x)H(x)\,.$
If $x\in Y$, then $H(x)=\mathrm{Id}$ and $P(x)=B(x)$, while if $x\not\in Y$, then the matrices $H(x)$ and $P(x)$ are (entrywise) positive. In both cases, $\det H(x)=1$ and $\det P(x)=\pm 1$.
The cocycle $(T,P)$ is called the positive perturbation of the cocycle $(T,B)$ and is an important intermediate step in our construction.
The following simple bound relies fundamentally on positivity and will be essential in the proof of Lemma 3.5 below.
Lemma 3.4.
For all $x\in X$ and $n\geq 0$, if $v$ is a nonnegative vector $v$, then
(3.22) $\|P^{(n)}(x)v\|\geq\theta(x)\|B^{(n)}(x)\|\,\|v\|\,.$
Proof.
Since $\cosh t\geq\sinh t\geq t$ for every $t\geq 0$, we have
(3.23) $H(x)\geq\theta(x)U\,,\quad\text{where}\quad U\coloneqq\begin{pmatrix}1&1\\ 1&1\end{pmatrix}$
and we are using the entrywise partial order on the set of matrices. On the other hand, since $\cosh t\geq 1$, we have $H(T^{i}x)\geq\mathrm{Id}$ for every $i$. Since product of nonnegative matrices is increasing with respect to the partial order, it follows that
(3.24) $P^{(n)}(x)$ $=B(T^{n-1}x)H(T^{n-1}x)\cdots B(Tx)H(Tx)B(x)H(x)$
(3.25) $\geq\theta(x)B(T^{n-1}x)\cdots B(Tx)B(x)U$
(3.26) $=\theta(x)B^{(n)}(x)U\,.$
Next, note that
(3.27) $Uv\geq\|v\|u\,,\quad\text{where}\quad u\coloneqq\begin{pmatrix}1\\ 1\end{pmatrix}\,.$
Let $w$ be a nonnegative unit (column-)vector such that $\|B^{(n)}(x)w\|=\|B^{(n)}(x)\|$. Note that $w\leq u$. Combining those observations,
(3.28) $P^{(n)}(x)$ $\geq\theta(x)B^{(n)}(x)Uv$
(3.29) $\geq\theta(x)\|v\|\,B^{(n)}(x)u$
(3.30) $\geq\theta(x)\|v\|\,B^{(n)}(x)w\,.$
Taking norms, we obtain the desired inequality. ∎
3.4. Avoiding cancellations
The next step establish the following lemma, which is a central component of our proof:
Lemma 3.5.
For all $k\in\mathbb{N}$, $x\in X$, $n\in\mathbb{Z}$, and $v$ a nonnegative unit vector, if
(3.31) $\delta_{k+1}\leq\mathsf{d}(x,Y)\leq\delta_{k}\quad\text{and}\quad m_{k}\leq n\leq m_{k+1}\,,$
then
(3.32) $\left|\frac{1}{n}\log\big{\|}P^{(n)}(x)v\big{\|}-\lambda_{0}\right|<\epsilon_{k}\,.$
Proof.
We can assume that $k>2$. Fix a point $x\in X$ such that $\delta_{k+1}\leq\mathsf{d}(x,Y)\leq\delta_{k}$. Let $y\in Y$ be such that $\mathsf{d}(x,y)=\mathsf{d}(x,Y)$. We begin the proof by establishing two simple but crucial inequalities, illustrated in Fig. 1. Since $\mathsf{d}(x,y)\leq\delta_{k}$, property (3.14) implies that
(3.33) $\mathsf{d}_{m_{k+1}}(x,y)\leq\delta_{k-1}\,.$
On the other hand, we claim that for all integers $i$ in the range $0\leq i\leq m_{k+1}$,
(3.34) $\mathsf{d}(T^{i}x,Y)\geq\delta_{k+2}\,.$
Indeed, if the inequality fails, then there exists $y^{\prime}\in Y$ such that $\mathsf{d}(T^{i}x,y^{\prime})<\delta_{k+2}$. In that case, property (3.14) gives $\mathsf{d}_{m_{k+3}}(T^{i}x,y^{\prime})<\delta_{k+1}$. In particular, since $|i|\leq m_{k+1}<m_{k+3}$, we have $\mathsf{d}(x,T^{-i}y^{\prime})<\delta_{k+1}$, contradicting the fact that $\mathsf{d}(x,Y)\geq\delta_{k+1}$.
[…]
Figure 1. Inequality (3.33) says that the points $x,Tx,\dots,T^{m_{k+1}}x$ never go too far away from $y,Ty,\dots,T^{m_{k+1}}y$ (respectively), while inequality (3.34) says that those points never get too close to $Y$.
Applying Lemma 3.2, we obtain a decomposition $n=r+qm_{k-1}+s$ where $q>0$, $|r|<m_{k-1}$, $|s|<m_{k-1}$, and and
(3.35) $\frac{1}{qm_{k-1}}\sum_{j=0}^{q-1}\log\|B^{(m_{k-1})}(T^{r+jm_{k-1}}y)\|>\lambda_{0}-\epsilon_{k}\,.$
Consider points $x_{j}\coloneqq T^{r+jm_{k-1}}x$. We factorize
(3.36) $P^{(n)}(x)=R_{2}P_{q-1}\cdots P_{1}P_{0}R_{1}\,,$
where $R_{1}\coloneqq P^{(r)}(x)$, $P_{j}\coloneqq P^{(m_{k-1})}(x_{j})$, and $R_{2}\coloneqq P^{(s)}(x_{r})$.
Let $\theta_{0}\coloneqq\phi(\delta_{k+2})$, where the function $\phi$ comes from Lemma 3.3. It follows from (3.34) that $\theta(x_{j})\geq\theta_{0}$. Therefore, by Lemma 3.4, for all nonnegative vectors $w$, we have
(3.37) $\|P_{j}w\|\geq\theta_{0}\|B_{j}\|\|w\|\,,\quad\text{where}\quad B_{j}\coloneqq B^{(m_{k-1})}(x_{j})\,.$
Applying this property recursively, we obtain
(3.38) $\|P_{q-1}\cdots P_{0}w\|\geq\theta_{0}^{q}\|B_{q-1}\|\cdots\|B_{0}\|\|w\|\,.$
By (3.16), we actually have $\theta_{0}=e^{-m_{k-2}}$, so
(3.39) $\theta_{0}^{q}=e^{-qm_{k-2}}>e^{-\epsilon_{k}qm_{k}}>e^{-\epsilon_{k}n}\,.$
Next, consider the points $y_{j}\coloneqq T^{r+jm_{k-1}}y$ and the corresponding matrix products $\tilde{B}_{j}\coloneqq B^{(m_{k-1})}(y_{j})$. By (3.33) and (3.15), we have
(3.40) $\|B_{j}\|>e^{-\epsilon_{k}m_{k-1}}\|\tilde{B}_{j}\|\quad\text{for each }j\in\ldbrack 0,q-1\rdbrack\,.$
In particular, $\|B_{q-1}\|\cdots\|B_{0}\|>e^{-\epsilon_{k}n}\|\tilde{B}_{q-1}\|\cdots\|\tilde{B}_{0}\|$. On the other hand, by (3.35), $\|\tilde{B}_{q-1}\|\cdots\|\tilde{B}_{0}\|>e^{(\lambda_{0}-\epsilon_{k})n}$. We combine the bounds above: given a nonnegative unit vector $v$, letting $w\coloneqq\frac{R_{1}v}{\|R_{1}v\|}$, we have
(3.41) $\|P^{(n)}(x)v\|$ $\geq\|R_{2}^{-1}\|^{-1}\,\|P_{q-1}\cdots P_{0}w\|\,\|R_{1}^{-1}\|^{-1}$
(3.42) $>e^{-Cm_{k-1}}\,\theta_{0}^{q}\,\|B_{q-1}\|\cdots\|B_{0}\|$
(3.43) $>e^{(\lambda_{0}-\epsilon_{k})n}\,.$
That is, $\frac{1}{n}\log\|P^{(n)}(x)v\|$ admits the lower bound $\lambda_{0}-\epsilon_{k}$, which confirms part of (3.32). Obtaining an upper bound uses different (and simpler) arguments, so we will prove it separately in Lemma 3.6. ∎
Lemma 3.6.
If $x\in X$, $k\in\mathbb{N}$, $n\geq m_{k}$, and $y\in Y$ are such that $\mathsf{d}_{n}(x,y)\leq\delta_{k-1}$, then $\frac{1}{n}\log\|P^{(n)}(x)\|<\lambda_{0}+\epsilon_{k}$.
Proof.
We can assume that $k>6$. Let $\ell\coloneqq m_{k-6}$. Perform Euclidean division:
(3.44) $n=q\ell+r\,,\quad q>0\,,\quad 0\leq r<\ell\,.$
For $j\in\ldbrack 0,q\rdbrack$, let $x_{j}\coloneqq T^{j\ell}(x)$. Consider the factorization
(3.45) $P^{(n)}(x)=P^{(r)}(x_{q})P^{(\ell)}(x_{q-1})\cdots P^{(\ell)}(x_{1})P^{(\ell)}(x_{0})\,.$
We will estimate the norm of each factor, starting with
(3.46) $\|P^{(r)}(x_{q})\|<e^{Cr}<e^{C\ell}=e^{Cm_{k-6}}<e^{\epsilon_{k}m_{k}}\leq e^{\epsilon_{k}n}\,.$
To bound the other factors, fix $j\in\ldbrack 0,q-1\rdbrack$. Note the telescopic sum
(3.47) $P^{(\ell)}(x_{j})-B^{(\ell)}(x_{j})=\sum_{i=0}^{\ell-1}P^{(\ell-i-1)}(T^{i+1}x_{j})\left[P(T^{i}x_{j})-B(T^{i}x_{j})\right]B^{(i)}(x_{j})\,.$
Taking norms,
(3.48) $\left\|P^{(\ell)}(x_{j})-B^{(\ell)}(x_{j})\right\|$ $\leq e^{C(\ell-1)}\sum_{i=0}^{\ell-1}\left\|P(T^{i}x_{j})-B(T^{i}x_{j})\right\|$
(3.49) $\leq e^{C\ell}\sum_{i=0}^{\ell-1}\left\|H(T^{i}x_{j})-\mathrm{Id}\right\|$
(3.50) $\leq e^{C\ell}\sum_{i=0}^{\ell-1}C\theta(T^{i}x_{j})\,.$
Since $\mathsf{d}(T^{i}x_{j})\leq\mathsf{d}_{n}(x,y)\leq\delta_{k-1}$, using (3.16) we get $\theta(T^{i}x_{j})\leq\phi(\delta_{k-1})=e^{-m_{k-5}}$. Since $\ell=m_{k-6}=\epsilon_{k}m_{k-5}$, the quantity in (3.50) is less than $1$ for large enough $k$. In particular,
(3.51) $\|P^{(\ell)}(x_{j})\|<1+\|B^{(\ell)}(x_{j})\|\leq 2\|B^{(\ell)}(x_{j})\|$
for $k$ large enough. Let $y_{j}\coloneqq T^{j\ell}(y)$. By (3.15), we have $\|B^{(\ell)}(x_{j})\|<2\|B^{(\ell)}(y_{j})\|$. On the other hand, limit (3.2) gives $\|B^{(\ell)}(y_{j})\|<e^{(\lambda_{0}+\epsilon_{k})\ell}$. It follows that $\|P^{(\ell)}(x_{j})\|<e^{(\lambda_{0}+\epsilon_{k})\ell}$. Using this bound together with (3.46) in the factorization (3.45), the outcome is $\|P^{(n)}(x)\|<e^{(\lambda_{0}+\epsilon_{k})n}$. Lemmas 3.6 and 3.5 are now proved. ∎
3.5. Invariant bundle
The goal of this subsection is showing that the restriction of the cocycle $(T,P)$ to the noncompact set $X\smallsetminus Y$ admits an invariant continuous field of directions (Lemma 3.8). The proof is independent of Lemma 3.5, and the key ingredient is the growth bound (3.17).
We will make use of Hilbert’s projective metric, which we now recall. Proof of the facts listed below can be found in [Bi, §2]. Consider the open positive cone $\mathbb{R}^{2}=\mathbb{R}\times\mathbb{R}$, i.e., the interior of the first quadrant in the Cartesian plane. For each $v=(v_{1},v_{2})\in\mathbb{R}^{2}$, the ray $\mathbb{R}v$ is denoted by $\vec{v}$ or $[v_{1}:v_{2}]$. Let $\mathbb{P}^{1}$ denote the set of such rays. We metrize this set with the Hilbert metric
(3.52) $\mathsf{d_{H}}\big{(}[v_{1}:v_{2}],[w_{1}:w_{2}]\big{)}\coloneqq\left|\log\frac{v_{1}w_{2}}{v_{2}w_{1}}\right|\,,$
or equivalently,
(3.53) $\mathsf{d_{H}}\big{(}[e^{t}:1],[e^{s}:1]\big{)}=|t-s|\,.$
Let $M$ be a nonnegative $2\times 2$ matrix such that $\mathrm{Ker}(M)\cap\mathbb{R}^{2}=\varnothing$. The projectivization of $M$ is the map $\vec{M}\colon\mathbb{P}^{1}\to\mathbb{P}^{1}$ defined by $\vec{M}(\vec{v})\coloneqq\overrightarrow{Mv}$. This map is Lipschitz with respect to the Hilbert metric; in fact, for all $\vec{v},\vec{w}\in\mathbb{P}^{1}$ we have
(3.54) $\mathsf{d_{H}}\big{(}\vec{M}(\vec{v}),\vec{M}(\vec{w})\big{)}\leq\left(\tanh\frac{\mathfrak{D}(M)}{4}\right)\,\mathsf{d_{H}}(\vec{v},\vec{w})\,,$
where $\mathfrak{D}(M)\in[0,\infty]$ denotes the diameter of the image $\vec{M}(\mathbb{P}^{1})$ with respect to the Hilbert metric. The Lipschitz constant $\tanh\frac{\mathfrak{D}(M)}{4}$ is at most $1$, and is strictly less than $1$ if the matrix has strictly positive entries. Actually,
(3.55) $M=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\quad\Rightarrow\quad\mathfrak{D}(M)=\left|\log\frac{ad}{bc}\right|\,,$
We also note that, if $M_{1}$ and $M_{2}$ are nonnegative matrices, then, as a direct consequence of bound (3.54),
(3.56) $\mathfrak{D}(M_{2}M_{1})\leq\left(\tanh\frac{\mathfrak{D}(M_{2})}{4}\right)\mathfrak{D}(M_{1})\,.$
Next, we establish the following “shrinking property” of the cocycle $(T,P)$:
Lemma 3.7.
For every $x\in X\smallsetminus Y$,
(3.57) $\lim_{n\to+\infty}\mathfrak{D}(P^{(n)}(T^{-n}x))=0\,.$
Proof.
Let $x\in X\smallsetminus Y$. Since $P(x)=B(x)H(x)$, we have $\mathfrak{D}(P(x))\leq\mathfrak{D}(H(x))$. On the other hand, by formula (3.55) and definition (3.20), $\mathfrak{D}(H(x))=-2\log\tanh\theta(x)$. Since the function $\theta$ is bounded above, we can write $\tanh\theta(x)\geq C^{-1}\theta(x)$ and therefore
(3.58) $\mathfrak{D}(P(x))\leq C-2\log\theta(x)\,.$
The following elementary bounds are valid for all $s>0$:
(3.59) $\log\tanh s=\log\frac{1-e^{-2s}}{1+e^{-2s}}<\log(1-e^{-2s})<-e^{-2s}\,.$
Note the following consequence:
(3.60) $\log\tanh\frac{\mathfrak{D}(P(x))}{4}\leq-C\theta(x)\,.$
Since $T$ is biLipschitz, we have, for all $j\geq 1$,
(3.61) $\mathsf{d}(T^{-j}x,Y)$ $\geq e^{-Cj}\mathsf{d}(x,Y)$
(3.62) $\geq e^{-\ell_{x}-Cj}\,,$
where $\ell_{x}\coloneqq\max\{0,-\log\mathsf{d}(x,Y)\}$. Using (3.19) and (3.17),
(3.63) $\theta(T^{-j}x)=\phi(\mathsf{d}(T^{-j}x,Y))$ $>\frac{1}{-\log\mathsf{d}(T^{-j}x,Y)}$
(3.64) $>\frac{1}{\ell_{x}+Cj}\quad\text{for all $j\geq 1$.}$
Consider the product $P^{(n)}(T^{-n}x)=P(T^{-1}x)\cdots P(T^{-(n-1)}x)P(T^{-n}x)$, where $n>0$. Applying (3.56) recursively,
(3.65) $\mathfrak{D}(P^{(n)}(T^{-n}x))\leq\mathfrak{D}(P(T^{-n}x))\prod_{j=1}^{n-1}\tanh\frac{\mathfrak{D}(P(T^{-j}x))}{4}$
Taking log’s and using (3.58) and (3.60), we obtain:
(3.66) $\log\mathfrak{D}(P^{(n)}(T^{-n}x))\leq\log\big{(}C-2\log\theta(T^{-n}x)\big{)}-C\sum_{j=1}^{n-1}\theta(T^{-j}x)\,.$
Now, using (3.64),
(3.67) $\log\mathfrak{D}(P^{(n)}(T^{-n}x))\leq\log\big{(}C+2\log(\ell_{x}+Cn)\big{)}-C\sum_{j=1}^{n-1}\frac{1}{\ell_{x}+Cj}\,.$
We bound the last sum a la Maclaurin–Cauchy: we have
(3.68) $\sum_{j=1}^{n-1}\frac{1}{\ell_{x}+Cj}>\int_{1}^{n}\frac{ds}{\ell_{x}+Cs}>\frac{1}{C}\log(\ell_{x}+Cn)\,,$
which diverges to +$\infty$+ ∞ as $n\to\infty$, and furthermore dominates the other term in the right-hand side of (3.67). This proves that $\mathfrak{D}(P^{(n)}(T^{-n}x))$ converges to 00. ∎
Lemma 3.8.
There exist continuous functions $u\colon X\smallsetminus Y\to\mathbb{R}^{2}$ and $f\colon X\smallsetminus Y\to\mathbb{R}$ such that for all $x\in X\smallsetminus Y$, we have $\|u(x)\|=1$ and
(3.69) $P(x)u(x)=e^{\lambda_{0}+f(x)}u(Tx)\,.$
Proof.
For each $x\in X\smallsetminus Y$ and $n>0$, let $\mathcal{K}^{n}_{x}$ be the closure of $\vec{P}^{(n)}(T^{-n}x)(\mathbb{P}^{1})$. Note that these sets are nested, that is, $\mathcal{K}^{n+1}_{x}\subseteq\mathcal{K}^{n}_{x}$. By Lemma 3.7, their Hilbert diameters converge to zero. It follows that $\bigcap_{n>0}\mathcal{K}^{n}_{x}$ contains a single point $\vec{u}(x)$ in $\mathbb{P}^{1}$, which can be expressed uniquely as $\vec{u}(x)=\overrightarrow{u(x)}$ where $u(x)$ is a positive unit vector. Since $\vec{P}(x)\mathcal{K}^{n}_{x}=\mathcal{K}^{n+1}_{Tx}$, it follows that $\vec{P}(x)\vec{u}(x)=\vec{u}(Tx)$. In particular, equation (3.69) uniquely determines the function $f$.
We are left to check continuity. Given $x\in X\smallsetminus Y$ and $\xi>0$, fix $n$ such that the diameter of $\mathcal{K}^{n}_{x}$ is less than $\xi$. Let $V$ denote the $\xi$-neighborhood of $\mathcal{K}^{n}_{x}$ with respect to the Hilbert metric. Since the matrix $P^{(n)}(T^{-n}x)$ has strictly positive entries that depend continuously on $x$, if $\tilde{x}$ is sufficiently close to $x$, then $\mathcal{K}^{n}_{\tilde{x}}\subseteq V$. Therefore $\vec{u}(\tilde{x})\in V$, and in particular, $\mathsf{d_{H}}(\vec{u}(\tilde{x}),\vec{u}(x))<2\xi$. This proves that the map $x\mapsto\vec{u}(x)\in\mathbb{P}^{1}$ is continuous with respect to the Hilbert metric. Equivalently, $x\mapsto u(x)\in\mathbb{R}^{2}$ is continuous with respect to the Euclidean metric. Lastly, $f(x)=-\lambda_{0}+\log\|P(x)u(x)\|$ is also a continuous function of $x$. ∎
3.6. The cohomological equation
The goal of this subsection is Lemma 3.12, which will show that the function $f$ from Lemma 3.8 is an “approximate coboundary.”
Lemma 3.9.
There exists a continuous function $\tau\colon X\smallsetminus Y\to\mathbb{R}$ such that
(3.70) $\mathsf{d}(x,Y)\leq\delta_{k}\quad\Rightarrow\quad\tau(x)\geq m_{k}\,,$
(3.71) $\lim_{x\to Y}|\tau(Tx)-\tau(x)|=0\,.$
Observe that $\tau(x)$ tends to infinity as $x$ approaches $Y$, but this convergence is slow, because $(\delta_{k})$ tends to zero much faster than $(m_{k})$ grows.
Proof.
Let $\psi\colon[\log 10,\infty)\to[1,\infty)$ be the piecewise affine homeomorphism with anchor points $\psi(\log 10)=1$ and $\psi(-\log\delta_{k})=m_{k}$. Define a continuous positive function on $X\smallsetminus Y$ by $\tau(x)\coloneqq\psi(-\log\mathsf{d}(x,Y))$. Property (3.70) is clearly satisfied.
Since our dynamics $T$ is a biLipschitz homeomorphism,
(3.72) $\left|-\log\mathsf{d}(Tx,Y)+\log\mathsf{d}(x,Y)\right|\leq C\,.$
So property (3.71) will follow once we check that the Lipschitz constant of the function $\psi$ restricted to the interval $[n,\infty)$ tends to zero as $n\to\infty$. The derivative of the piecewise linear function $\psi$ between two consecutive anchor points is:
[…]
(3.74) $<\frac{1}{k}\,,$
so the function $\psi$ behaves appropriately. ∎
Recall that the function $f$ was defined in Lemma 3.8. Note that, despite its domain being noncompact, $f$ is bounded. Let $\|f\|_{\infty}\coloneqq\sup|f|$.
For each $t\geq 0$ and $x\in X\smallsetminus Y$, we define the interpolated Birkhoff sum $S_{t}(x)$ as follows: $S_{0}(x)\coloneqq 0$ and, if $t>0$,
(3.75) $S_{t}(x)\coloneqq f(x)+f(Tx)+\cdots+f(T^{\lfloor t\rfloor-1}x)+(t-\lfloor t\rfloor)f(T^{\lfloor t\rfloor}x)\,.$
Also, define the auxiliary function
(3.76) $Z_{t}(x)\coloneqq\sum_{i=0}^{\infty}\left(1-\frac{i+1}{t}\right)^{+}f(T^{i}x)\,.$
Lemma 3.10.
The function $(t,x)\mapsto Z_{t}(x)$ is continuous on ×$\mathbb{R}\times(X\smallsetminus Y)$. It solves the following cohomological equation,
(3.77) $\frac{S_{t}(x)}{t}=f(x)+Z_{t}(Tx)-Z_{t}(x)\,,$
for all $t>0$. Furthermore, for all $t,s>0$,
(3.78) $|Z_{s}(x)-Z_{t}(x)|\leq\|f\|_{\infty}\,|s-t|\,.$
We remark that (3.77) is basically known, at least for integer $t$ (see [MP]), while property (3.78) seems to be new.
Proof.
To check continuity, note the terms in the sum (3.76) vanish for $i\geq t-1$, and are continuous with respect to $x$ and $t$.
Note that (3.75) can be rewritten as:
(3.79) $S_{t}(x)=\sum_{i=0}^{\infty}\left[(t-i)^{+}-(t-i-1)^{+}\right]f(T^{i}x)\,.$
So:
(3.80) $\frac{S_{t}(x)}{t}=\underbrace{\sum_{i=0}^{\infty}\left(1-\frac{i}{t}\right)^{+}f(T^{i}x)}_{f(x)-Z_{t}(Tx)}-\underbrace{\sum_{i=0}^{\infty}\left(1-\frac{i+1}{t}\right)^{+}f(T^{i}x)}_{Z_{t}(x)}\,,$
which is equation (3.77).
Since the coefficient of $f(T^{i}x)$ in the series (3.75) is increasing with respect to $t$, we have $|Z_{s}(x)-Z_{t}(x)|\leq\|f\|_{\infty}|h(s)-h(t)|$, where
(3.81) $h(t)\coloneqq\sum_{i=0}^{\infty}\left(1-\frac{i+1}{t}\right)^{+}\,.$
This is a continuous function. If $t>0$ is not an integer, then $h$ is differentiable at $t$ with
(3.82) $h^{\prime}(t)=\sum_{i=0}^{\lfloor t\rfloor-1}\frac{i+1}{t^{2}}=\frac{\lfloor t\rfloor\,(\lfloor t\rfloor+1)}{2t^{2}}\in[0,1]\,.$
So $h$ has a Lipschitz contant $1$, and inequality (3.78) follows. ∎
Lemma 3.11.
We have
(3.83) $\lim_{x\to Y}\frac{S_{\tau(x)}(x)}{\tau(x)}=0\,.$
Proof.
Given $x\in X\smallsetminus Y$, let $k$ be such that $\delta_{k+1}\leq d(x,Y)\leq\delta_{k}$. Then (3.70) implies that $m_{k}\leq\tau(x)\leq m_{k+1}$. Let $n\coloneqq\lfloor\tau(x)\rfloor$, so $m_{k}\leq n\leq m_{k+1}$, that is, conditions (3.31) are met. Now write $\tau(x)=n+r$, where $10\leq r<$ 1. By relations (3.75) and (3.69),
(3.84) $\frac{S_{\tau(x)}(x)}{\tau(x)}$ $=-\lambda_{0}+\frac{1}{n+r}\log\|P^{(n)}(x)u(x)\|+\frac{r}{n+r}\log\|P(T^{n}x)u(T^{n}x)\|$
(3.85) $=-\lambda_{0}+\frac{1}{n}\log\|P^{(n)}(x)u(x)\|+O(\tfrac{1}{n})\,.$
If $x$ is close to $Y$, then $k$ and $n$ are large, and Lemma 3.5 ensures that this quantity is small. ∎
Lemma 3.12 (Approximate coboundary).
There exist two continuous functions $g\colon X\smallsetminus Y\to\mathbb{R}$ and $r\colon X\to\mathbb{R}$ such that
(3.86) $f(x)=g(Tx)-g(x)+r(x)\quad\text{for all }x\in X\smallsetminus Y$
and $r$ vanishes on $Y$.
Proof.
We let $g(x)\coloneqq-Z_{\tau(x)}(x)$. This is a continuous function on $X\smallsetminus Y$ (see Lemma 3.10). Now define $r$ on $X\smallsetminus Y$ by (3.86). Then we have
(3.87) $|r(x)|$ $=\left|f(x)+g(x)-g(Tx)\right|$
(3.88) $=\left|f(x)-Z_{\tau(x)}(x)+Z_{\tau(Tx)}(Tx)\right|$
(3.89) $\leq\left|f(x)-Z_{\tau(x)}(x)+Z_{\tau(x)}(Tx)\right|+\big{|}Z_{\tau(Tx)}(Tx)-Z_{\tau(x)}(Tx)\big{|}$
Therefore
(3.90) $|r(x)|$ $\leq\left|\frac{S_{\tau(x)}(x)}{\tau(x)}\right|+\|f\|_{\infty}\left|\tau(Tx)-\tau(x)\right|$ (by (3.77) and (3.78))
(3.91) $\to 0\text{ as $x\to Y$}$ (by (3.83) and (3.71)).
This allows us to define $r$ as 00 on $Y$ and obtain a continuous function on $X$. ∎
3.7. The final perturbation
For each $x\in X\smallsetminus Y$, let $S(x)$ be the matrix uniquely defined by the following conditions:
(3.92) $S(x)u(x)=e^{r(x)}u(x)\quad\text{and}\quad S(x)v=e^{-r(x)}v\text{ if }v\perp u(x)\,.$
Since $u(x)$ and $r(x)$ depend continuously on $x$, so does $S(x)$. Furthermore, $S(x)$ is a symmetric matrix with eigenvalues $e^{\pm r(x)}$. It follows that
(3.93) $\|S(x)-\mathrm{Id}\|=e^{|r(x)|}-1\,.$
Since the continuous function $r$ vanishes on $Y$, defining $S(x)$ as the identity matrix for all $x\in Y$, we obtain a continuous map $S\colon X\to\mathrm{SL}(2,\mathbb{R})$. Finally, define $A\colon X\to\mathrm{SL}(2,\mathbb{R})$ by
(3.94) $A(x)=P(x)S(x)\,.$
Note that
(3.95) $A(x)u(x)=e^{\lambda_{0}+g(Tx)-g(x)}u(Tx)\,,\quad\text{for all $x\in X\smallsetminus Y$.}$
The cocycle $(T,A)$ is not uniformly hyperbolic, since its restriction to the invariant set $Y$ is still the Walters cocycle. Let us show that the cocycle is monochromatic:
Lemma 3.13.
$\lambda(T,A,\mu)=\lambda_{0}$ for every ergodic measure $\mu$.
Proof.
Let $\mu$ be a $T$-invariant ergodic Borel probability measure on $X$. If $\mu(Y)=1$, then $\mu$ equals $\nu$, the unique invariant measure for the Veech map, so the Lyapunov exponent is $\lambda_{0}$ by definition. So we can assume that $\mu(Y)=0$.
For every $x\in X\smallsetminus Y$ and $n\geq 0$, iteration of (3.95) gives
(3.96) $A^{(n)}(x)u(x)=e^{n\lambda_{0}+g(T^{n}x)-g(x)}u(T^{n}(x))\,.$
By the Poincaré recurrence theorem, for $\mu$-almost every $x$ there exists a sequence $n_{i}\to+\infty$ such that $T^{n_{i}}x\to x$. We can assume that $x\not\in Y$, so
(3.97) $\frac{1}{n_{i}}\log\|A^{(n_{i})}(x)u(x)\|=\lambda_{0}+\frac{g(T^{n}x)-g(x)}{n_{i}}\to\lambda_{0}\quad\text{as }i\to\infty\,,$
since $g$ is continuous on the open set $X\smallsetminus Y$. On the other hand, by Oseledets theorem, for $\mu$-almost every $x$ and every non-zero vector $v$ we have
(3.98) $\frac{1}{n}\log\|A^{(n_{i})}(x)v\|\to\pm\lambda(T,A,\mu)\quad\text{as }n\to\infty\,.$
It follows that $\lambda(T,A,\mu)=\lambda_{0}$. ∎
This concludes the proof of Theorem 2.6. Observe that it is not necessarily the case that matrices $A(x)$ are positive (or even nonnegative) for $x\in X\smallsetminus Y$. Nevertheless, the cocycle restricted to $X\smallsetminus Y$ admits a continuous field of lines such along which the log of the rate of expansion is cohomologous to a constant. In particular, the restriction of the cocycle to any invariant compact set $Z\subset X\smallsetminus Y$ is uniformly hyperbolic.
4. Construction of a symbolic Veech map
In this section we provide a proof of Theorem 2.2. We use a variation of a construction explained to us by Scott Schmieding.
4.1. Definition of the shift space
Consider the following alphabet
[…]
We fix an integer sequence $(m_{k})$ such that
(4.2) $m_{k}\geq 2\quad\text{and}\quad\sum_{k=1}^{\infty}\frac{1}{m_{k}}<\infty\,.$
Then we define recursively a sequence of words $e_{1},e_{2},\dots$ as follows:
[…]
Let $Y$ be the set of doubly-infinite sequences $\omega\in\mathsf{A}^{\mathbb{Z}}$ such that for every $n>0$, the word $\omega_{\ldbrack-n,n\rdbrack}$ is a subword of some $e_{k}$. Equipping $\mathsf{A}^{\mathbb{Z}}$ with the product topology, the subset $Y$ is closed and shift-invariant. The fact that $Y$ is nonempty follows from the observation that for every $k$, the word $e_{k}$ can be extended both to the left and to the right to form the word $e_{k+1}$. Let $T\colon Y\to Y$ be the restriction of the shift transformation. We will show:
Theorem 4.1.
The map $T$ defined above is Veech, and has zero topological entropy.
The theorem above obviously implies Theorem 2.2. We proceed with the proof, which is divided into several steps.
4.2. Entropy and minimality
Let $\mathcal{L}=\mathcal{L}(Y)$ be the language of $Y$, that is, the set of all words that occur in elements of $Y$. Equivalently, $w\in\mathcal{L}$ if and only if $w$ is a subword of some $e_{k}$.
For each $k>0$, the four words
(4.5) $e_{k}\,,\quad\overline{e_{k}}\,,\quad e_{k}\,0\,,\quad\text{and}\quad\overline{e_{k}}\,0$
will be called $k$-elementary words.
Lemma 4.2.
For each $k>0$, let $\Sigma_{k}\subseteq\mathsf{A}^{\mathbb{Z}}$ be the shift space formed by all bi-infinite words obtained by concatenation of $k$-elementary words. Then $\Sigma_{k}\supseteq Y$.
Before proving the Lemma, it is convenient to have an alternative description of the shift $\Sigma_{k}$. We use the terminology from [LM, §3.1]. Let $\mathcal{G}_{k}$ be a directed graph consisting of four loops based on a common vertex $\bullet$ (that is, a “four-leaf clover”), two of which with length $|e_{k}|$ and the other two with length $|e_{k}|+1$. Label the edges in such a way that the label sequences corresponding to the loops are the four $k$-elementary words (4.5). For example, the graph $\mathcal{G}_{1}$ is depicted in Fig. 2. Then $\Sigma_{k}$ is the set of label sequences of infinite walks on $\mathcal{G}_{k}$. That is, $\Sigma_{k}$ is the sofic shift presented by $\mathcal{G}_{k}$. The language $\mathcal{L}(\Sigma_{k})$ of the shift $\Sigma_{k}$ is formed by all label sequences of finite walks in the graph.
[…]
Proof of Lemma 4.2.
We claim that every $\ell$-elementary word with $\ell\geq k$ is the label sequence of a closed walk in the graph $\mathcal{G}_{\ell}$ starting and ending at the center vertex $\bullet$. The claim is proved by induction on $\ell$, using the recursive relation (4.4). As a consequence of the claim, $e_{\ell}\in\mathcal{L}(\Sigma_{k})$ for every $\ell$.
Now if $w$ is any word in $\mathcal{L}(Y)$, by definition $w$ is a subword of some $e_{\ell}$. In particular, $w\in\mathcal{L}(\Sigma_{k})$. We have proved that $\mathcal{L}(Y)\subseteq\mathcal{L}(\Sigma_{k})$. It follows (see [LM, p. 10]) that $Y\subseteq\Sigma_{k}$. ∎
Lemma 4.3.
The map $T$ has zero topological entropy.
Proof.
It is clear that the shift $\Sigma_{k}$ in Lemma 4.2 has entropy $h(\Sigma_{k})\leq\frac{\log 4}{|e_{k}|}$. Therefore, $h(Y)=0$. ∎
Lemma 4.4.
$T$ and $T^{2}$ are minimal.
Obviously, minimality of $T^{2}$ implies minimality of $T$, but for the sake of clarity we consider $T$ first.
Proof.
Recall that a homeomorphism of a compact metric space is minimal if and only if for every nonempty open subset $U$, there exists $n_{0}>0$ such that every segment of orbit of length $n_{0}$ intersects $U$ (see [FH, Prop. 1.6.25]). Therefore, in the case under consideration, proving that $T$ is minimal is tantamount to showing that given any word $v\in\mathcal{L}$, every sufficiently long word $w\in\mathcal{L}$ contains $v$ as a subword. Let us check this property. Given $v$, there exists $k>0$ such that $v$ is a subword of either $e_{k}$ or $\overline{e_{k}}$. Now suppose that $w\in\mathcal{L}$ has length $|w|>2|e_{k}|$. Then $w$ corresponds to a walk of length $>2|e_{k}|$ on the graph $\mathcal{G}_{k}$. This walk must traverse one of the four loops. That is, $w$ contains a subword which is $k$-elementary. In particular, $w$ has $v$ as a subword. We conclude that $T$ is minimal.
In order to prove that $T^{2}$ is minimal, it is sufficient to check that given any word $v\in\mathcal{L}$, every sufficiently long word $w\in\mathcal{L}$ contains $v$ as a subword in both odd and even positions. The argument is similar: Given $v\in\mathcal{L}$, take $k$ such that $v$ is a subword of either $e_{k}$ or $\overline{e_{k}}$. Now let $w\in\mathcal{L}$ have length $|w|>2|e_{k+1}|$. Then $w$ contains a subword $f$ which is one of the four $k+1$-elementary words. In any case, $e_{k}$ appears in $f$ in both odd and even positions. In particular, $v$ occurs as a subword of $w$ in both odd and even positions. This shows that $T^{2}$ is minimal. ∎
4.3. Unique ergodicity of $T$
Given two words $v$ and $w$ in $\mathsf{A}^{*}$, where $w$ is nonempty, let $|w|_{v}$ denote the number of occurrences of $v$ as a subword of $w$, that is, the number of possible ways of expressing $w$ as a concatenation $uvu^{\prime}$. Note that
(4.6) $|w|_{v}\leq\max\left(0,|w|-|v|+1\right)\leq|w|\,.$
Lemma 4.5.
For all words $v,w_{1},\dots,w_{p}$ in $\mathsf{A}^{*}$, with $v\neq\varnothing$, we have
(4.7) $\sum_{i=1}^{p}|w_{i}|_{v}\leq|w_{1}\cdots w_{p}|_{v}\leq\sum_{i=1}^{p}|w_{i}|_{v}\ +\ (p-1)(|v|-1)\,.$
Proof.
We claim that $|w_{1}w_{2}|_{v}$ is at least $|w_{1}|_{v}+|w_{2}|_{v}$ and at most $|w_{1}|_{v}+|w_{2}|_{v}+|v|-1$. Indeed, the lower bound corresponds to “internal” occurrences of $v$ as a subword of either $w_{1}$ or $w_{2}$, while the upper bound takes into account possible “concatenation” occurrences. So the Lemma is true in the case $p=2$ (and tautological in the case $p=1$). An induction on $p$ concludes the proof. ∎
For every nonempty word $v$ in the language $\mathcal{L}$, we will associated a “limit frequency” $\Phi(v)$, as follows:
Lemma 4.6.
For any nonempty $v\in\mathcal{L}$, the sequences $\frac{|e_{k}|_{v}}{|e_{k}|}$ and $\frac{|\overline{e_{k}}|_{v}}{|e_{k}|}$ converge to common limit $\Phi(v)$, which is strictly positive.
Proof.
Fix a nonempty word $v\in\mathcal{L}$. Applying Lemma 4.5 to the recursive formula (4.4), we obtain
(4.8) $(m_{k}+1)(|e_{k}|_{v}+|\overline{e_{k}}|_{v})\leq|e_{k+1}|_{v}\leq(m_{k}+1)(|e_{k}|_{v}+|\overline{e_{k}}|_{v})+(2m_{k}+3)(|v|-1)\,.$
In particular,
(4.9) $|e_{k+1}|_{v}=(m_{k}+1)(|e_{k}|_{v}+|\overline{e_{k}}|_{v})+O(m_{k})\,.$
Similarly,
(4.10) $|\overline{e_{k+1}}|_{v}=(m_{k}+1)(|e_{k}|_{v}+|\overline{e_{k}}|_{v})+O(m_{k})\,.$
Let $S_{k}\coloneqq|e_{k}|_{v}+|\overline{e_{k}}|_{v}$ and $D_{k}\coloneqq|e_{k}|_{v}-|\overline{e_{k}}|_{v}$. Then
(4.11) $S_{k+1}$ $=2(m_{k}+1)S_{k}+O(m_{k})\quad\text{and}$
(4.12) $D_{k+1}$ $=O(m_{k})\,.$
There exists $k_{0}$ such that $S_{k_{0}}>0$. If $k\geq k_{0}$, then $S_{k}>2^{k-k_{0}}S_{k_{0}}$ and
(4.13) $\frac{S_{k+1}}{S_{k}}=2(m_{k}+1)+O\left(\frac{m_{k}}{S_{k}}\right)$ $=2(m_{k}+1)\left(1+O\left(\frac{1}{S_{k}}\right)\right)$
(4.14) $=2(m_{k}+1)\left(1+O(2^{-k})\right)\,.$
On the other hand, by the recursion (4.4),
(4.15) $|e_{k+1}|=2(m_{k}+1)|e_{k}|+2\,.$
So $|e_{k}|\geq 2^{k}$ and
(4.16) $\frac{|e_{k+1}|}{|e_{k}|}=2(m_{k}+1)\left(1+O(2^{-k})\right)\,.$
Using (4.14) we obtain that for all $k\geq k_{0}$,
(4.17) $\frac{S_{k+1}}{|e_{k+1}|}=\frac{S_{k}}{|e_{k}|}\,r_{k}\,,\quad\text{where }r_{k}>0\text{ and }r_{k}=1+O(2^{-k})\,.$
It follows that $\frac{S_{k}}{|e_{k}|}=\frac{S_{k_{0}}}{|e_{k_{0}}|}\prod_{j=k_{0}}^{k}r_{j}$ converges as $k\to\infty$ to a finite non-zero limit, which we call $2\Phi(v)$.
Next, by (4.12) and (4.15), we have $\frac{D_{k+1}}{|e_{k+1}|}=O\left(\frac{1}{|e_{k}|}\right)=o(1)$. Therefore the two sequences
(4.18) $\frac{|e_{k}|_{v}}{|e_{k}|}=\frac{1}{2}\left(\frac{S_{k}}{|e_{k}|}+\frac{D_{k}}{|e_{k}|}\right)\quad\text{and}\quad\frac{|\overline{e_{k}}|_{v}}{|e_{k}|}=\frac{1}{2}\left(\frac{S_{k}}{|e_{k}|}-\frac{D_{k}}{|e_{k}|}\right)$
converge to the same positive number $\Phi(v)$, as we wanted to show. ∎
The next lemma builds on Lemma 4.6:
Lemma 4.7.
For every nonempty word $v\in\mathcal{L}$ and every $\varepsilon>0$, there exists $n>0$ such that if $w\in\mathcal{L}$ and $|w|>n$, then $\frac{|w|_{v}}{|w|}=\Phi(v)+O(\varepsilon)$.
Proof.
Fix a nonempty word $s\in\mathcal{L}$ and a positive number $\varepsilon$. Using Lemma 4.6, we see that if $k$ is large enough, we have
(4.19) $\left|\frac{|u|_{v}}{|f|}-\Phi(v)\right|<\varepsilon\quad\text{for each $k$-elementary word $f$.}$
Fix such a large enough $k$ with $|v|<\varepsilon|e_{k}|$. Then choose an integer $n>\varepsilon^{-1}(|e_{k}|+1)$, and let $w$ be any word in $\mathcal{L}$ with length at least $n$. As a consequence of Lemma 4.2, the word $w$ admits a factorization
(4.20) $w=f_{1}f_{2}\cdots f_{p-1}f_{p}$
where the inner factors $f_{2},\dots,f_{p-1}$ are $k$-elementary words, and the extremal factors $f_{1}$ and $f_{p}$ have length at most $|e_{k}|$. Note that $|w|\geq(p-2)|e_{k}|$. Applying Lemma 4.5 to the factorization (4.20), we obtain
(4.21) $\sum_{i=2}^{p-1}|f_{i}|_{v}\leq|w|_{v}<\sum_{i=1}^{p}|f_{p}|_{v}+p|v|\,.$
The difference between the upper and the lower bound is
(4.22) $|f_{0}|+|f_{1}|+p|v|<2|e_{k}|+p\varepsilon|e_{k}|\leq 2|e_{k}|+\varepsilon(|w|-2|e_{k}|)=O(\varepsilon|w|).$
For each $i$ between $2$ and $p-1$, by (4.19) we have $|f_{i}|_{v}=\Phi(v)|f_{i}|+O(\varepsilon|f_{i}|)$. Also,
(4.23) $\sum_{i=2}^{p-1}|f_{i}|=|w|-|f_{0}|-|f_{1}|=|w|+O(\varepsilon|w|)\,.$
It follows that
(4.24) $|w|_{v}=\sum_{i=2}^{p-1}|f_{i}|_{v}+O(\varepsilon|w|)=\Phi(v)|w|+O(\varepsilon|w|)\,,$
as we wanted to show. ∎
Lemma 4.8.
$T$ is uniquely ergodic.
Proof.
Given a nonempty word $v$ in $\mathcal{L}$, consider the cylinder [${}_{0}[v]$ consisting of all $\omega\in Y$ such that $\omega_{\ldbrack 0,|v|-1\rdbrack}=s$. It follows from Lemma 4.7 that the Birkhoff averages of the characteristic function of this cylinder converge uniformly to a constant, namely $\Phi(v)$. The same is true for more general cylinders [${}_{n}[v]\coloneqq T^{-n}({}_{0}[v])$. The set of linear combinations of characteristic functions of cylinders is dense in $C^{0}(Y,\mathbb{R})$. Therefore (see [KH, Proposition 4.1.15]), $T$ is uniquely ergodic. ∎
4.4. Non-unique ergodicity of $T^{2}$
Let us define two auxiliary functions $\varphi$ , $\psi$ on $Y$ as follows:
(4.25) $\varphi(\omega)$$\coloneqq\begin{cases}\phantom{-}1&\text{if }\omega_{0}=\mathord{\uparrow}\,,\\ -1&\text{if }\omega_{0}=\mathord{\downarrow}\,,\\ \phantom{-}0&\text{if }\omega_{0}=0\,,\end{cases}$
(4.26) $\psi$ $\coloneqq\varphi-\varphi\circ T\,.$
[…]
(4.27) $\varphi(\omega)=\omega_{0}\quad\text{and}\quad\psi(\omega)=\omega_{0}-\omega_{1}\,.$
We define another auxiliary function $\theta\colon\mathcal{L}\to\mathbb{R}$ as follows: $\theta(\varnothing)=0$ and, if $w\neq\varnothing$,
(4.28) $\theta(w)\coloneqq\sum_{i=0}^{|w|-1}(-1)^{i}\varphi(T^{i}(\omega))\,,$
where $\omega$ is any element of the cylinder [${}_{0}[w]$. So, in terms of the letter identifications $\mathord{\uparrow}=1$ and $\mathord{\downarrow}=-1$, if the word $w$ is spelt $a_{0}\cdots a_{n-1}$, then $\theta(w)=\sum_{i=0}^{n-1}(-1)^{i}a_{i}$. The function $\theta$ has the following properties:
(4.29) $\theta(\overline{w})=-\theta(w)\,,\quad\theta(wv)=\theta(w)+(-1)^{|w|}\theta(v)$
Lemma 4.9.
The sequence $\frac{\theta(e_{k})}{|e_{k}|}$ converges as $k\to\infty$ to a number $c>0$.
Proof.
Using the definitions (4.3), (4.4) and the properties (4.29), we obtain:
(4.30) $\theta(e_{1})=2\quad\text{and}\quad\theta(e_{k+1})=2(m_{k}-1)\theta(e_{k})\,.$
Using the recursive relation (4.15), we obtain
(4.31) $\frac{\theta(e_{k+1})}{|e_{k+1}|}=\frac{\theta(e_{k})}{|e_{k}|}\,r_{k}\quad\text{where}\quad r_{k}=\frac{m_{k}-1}{m_{k}+1+|e_{k}|^{-1}}\,.$
Since $m_{k}\geq 2$ and $|e_{k}|\geq 2$, we have
(4.32) $10<\frac{m_{k}-1}{m_{k}+1.5}\leq r_{k}<$ 1
and
(4.33) $\sum_{k=1}^{\infty}(1-r_{k})\leq\sum_{k=1}^{\infty}\frac{2.5}{m_{k}+1.5}<\infty\,,$
by the convergence condition in (4.2). If follows that $c\coloneqq\lim\frac{\theta(e_{k})}{|e_{k}|}$ exists and is positive. ∎
Lemma 4.10.
$T^{2}$ is not uniquely ergodic.
Proof.
[…]
(4.34) $\frac{1}{n}\sum_{i=0}^{n-1}\psi(T^{2i}(\omega))=\frac{2\theta(\omega_{\ldbrack 0,|e_{k}|-1\rdbrack})}{|e_{k}|}=\frac{2\theta(e_{k})}{|e_{k}|}\,.$
By Lemma 4.9, this quantity is close to $2c\neq 0$ if $k$ is large enough. So $T^{2}$ cannot be uniquely ergodic. ∎
The combination of Lemmas 4.3, 4.4, 4.8, and 4.10 yields Theorem 4.1.
4.5. Appendix: An explicit Walters cocycle
We have explicitly constructed a Veech map $T$ in Theorem 4.1. Then Theorem 2.4 guarantees the existence of a Walters cocycle with base $T$. However, the proof of the latter Theorem is slightly indirect. For the sake of completeness, let us provide an explicit example of Walters cocycle with base $T$.
Theorem 4.11.
Let $T\colon Y\to Y$ be the Veech map constructed above. Let $B\colon Y\to\mathrm{SL}(2,\mathbb{R})$ be defined by
(4.35) $B(\omega)=\begin{pmatrix}0&e^{\varphi(\omega)}\\ e^{-\varphi(\omega)}&0\end{pmatrix}\,,$
where $\varphi$ is as in (4.25). Then $(T,B)$ is a Walters cocycle, and its Lyapunov exponent with respect to the unique invariant measure equals the number $c$ from Lemma 4.9.
The proof needs the following observation:
Lemma 4.12.
Let $\omega\in Y$ be such that the limits
(4.36) $\lim_{n\to\infty}\frac{1}{n}\left|\sum_{j=0}^{n-1}(-1)^{j}\varphi(T^{j}(\omega))\right|\quad\text{and}\quad\lim_{n\to\infty}\frac{1}{n}\left|\sum_{j=-n}^{-1}(-1)^{j}\varphi(T^{j}(\omega))\right|$
both exist. Then at least one of them equals $c$.
Proof.
Let $\omega\in Y$, $\varepsilon>0$, and $n_{0}>0$ be arbitrary. We will show that there exists $n>n_{0}$ such that either $\frac{1}{n}\left|\sum_{j=0}^{n-1}(-1)^{j}\varphi(T^{j}(\omega))\right|$ or $\frac{1}{n}\left|\sum_{j=-n}^{-1}(-1)^{j}\varphi(T^{j}(\omega))\right|$ is $c+O(\varepsilon)$. Once we do that, the Lemma will be proved.
Choose $k>0$ satisfying the following conditions:
(4.37) $|e_{k+1}|>2n_{0}\,,\quad m_{k}>\frac{1}{\varepsilon}\,,\quad\text{and}\quad\left|\frac{\theta(e_{k})}{|e_{k}|}-c\right|<\varepsilon\,.$
By Lemma 4.2, $\omega\in\Sigma_{k+1}$, that is, the bi-infinite word $\omega$ can be written as a concatenation $\cdots f_{-1}f_{0}f_{1}\cdots$ of $k+1$-elementary words $f_{i}$, where the zeroth position is somewhere in $f_{0}$. Write $f_{0}=\omega_{\ldbrack-p,q-1\rdbrack}$, where $p\geq 0$ and $q\geq 1$. Let
(4.38) $n\coloneqq\max(p,q)\quad\text{and}\quad w\coloneqq\begin{cases}\omega_{\ldbrack-p,-1\rdbrack}&\text{if }n=p\,,\\ \omega_{\ldbrack 0,q-1\rdbrack}&\text{if }n=q\,.\end{cases}$
Since $p+q=|f_{0}|$, we have
(4.39) $|w|=n\geq\frac{|f_{0}|}{2}\geq\frac{|e_{k+1}|}{2}>n_{0}\,.$
To complete the proof, we will check that $\frac{|\theta(w)|}{|w|}=c+O(\varepsilon)$.
Consider the case $n=p$. Then $w$ is a prefix of the $k+1$-elementary word $f_{0}$. Since $\theta(\overline{w})=-\theta(w)$, we can assume that $f_{0}$ is either $e_{k+1}$ or $e_{k+1}0$. Since $|w|\geq\frac{1}{2}|f_{0}|$, it follows from the recursive definition (4.4) that, by either adding or removing at most $|e_{k}|+2$ letters at the end of the word $w$, we can obtain a word of the form
(4.40) $\tilde{w}=e_{k}^{m_{k}}\,0\,e_{k}\,\overline{e_{k}}^{r}\,,\quad\text{where }r\geq 0\,.$
Then
(4.41) $|w|=(m_{k}+r)|e_{k}|+O(|e_{k}|)\,.$
Also, by properties (4.29) we have $\theta(\tilde{w})=(m_{k}-1+r)\theta(e_{k})$ and $\big{|}\theta(w)-\theta(\tilde{w})\big{|}\leq|e_{k}|+2$, so
(4.42) $\theta(w)=(m_{k}+r)\theta(e_{k})+O(|e_{k}|)\,.$
Then, using a standard Calculus estimate,
(4.43) $\frac{\theta(w)}{|w|}=\frac{\theta(e_{k})}{|e_{k}|}+O\left(\frac{1}{m_{k}+r}\right)=c+O(\varepsilon)\,,$
as we wanted to show. This completes the argument in the case $n=p$. The case $n=q$ is handled in a similar manner. ∎
Proof of Theorem 4.11.
[…]
5. Questions
In this paper, we showed that the hypothesis of Hölder regularity cannot be dropped from the results of [DG, Gu]. Nevertheless, we do not know the exact regularity threshold for the validity of such theorems.
Question 5.1.
What is the optimal modulus of regularity for a map $A$ as in Theorem 1.1?
We did not try to estimate the modulus of regularity of the maps $A$ arising from our construction.
The monochromatic cocycles constructed in this paper take values in the disconnected Lie group $\mathrm{SL}(2,\mathbb{R})$. If we want examples taking values in a connected Lie group, we can use the fact that $\mathrm{SL}(2,\mathbb{R})$ embeds into $\mathrm{SL}(d,\mathbb{R})$, for any $d>2$.
We can also construct some examples taking values in $\mathrm{SL}(2,\mathbb{R})$, for instance:
Proposition 5.2.
If $T=\sigma_{4}\colon 4^{\mathbb{Z}}\to 4^{\mathbb{Z}}$ is the full shift on four symbols, then for any $\lambda_{0}>0$ we can find a map $A\colon 4^{\mathbb{Z}}\to\mathrm{SL}(2,\mathbb{R})$ satisfying the conclusions of Theorem 1.1.
Proof.
Let $\sigma_{2}$ be the full shift on two symbols. By Theorem 1.1, there exists $B\colon 2^{\mathbb{Z}}\to\mathrm{SL}(2,\mathbb{R})$ such that $\lambda(\sigma_{2},B,\mu)=\lambda_{0}/2$ for every $\sigma_{2}$-invariant ergodic measure $\mu$ and the cocycle $(\sigma_{2},B)$ is not uniformly hyperbolic. The proof of the theorem shows that $B$ can be taken with constant determinant $-1$. In particular, $B^{(2)}(x)=B(\sigma_{2}(x))B(x)$ has determinant $1$. Note that $\sigma_{2}$ is a square root of $\sigma_{4}$; more precisely, there exists a homeomorphism $h\colon 4^{\mathbb{Z}}\to 2^{\mathbb{Z}}$ such that $h\circ\sigma_{4}=\sigma_{2}^{2}\circ h$. Therefore $A\coloneqq B^{(2)}\circ h$ has the required properties. ∎
Question 5.3.
If $T=\sigma_{2}$ is the full shift on two symbols and $\lambda_{0}>0$, can we find a map $A\colon 2^{\mathbb{Z}}\to\mathrm{SL}(2,\mathbb{R})$ satisfying the conclusions of Theorem 1.1?
Acknowledgements. I thank Omri Sarig and Scott Schmieding for invaluable help with the construction of a Veech map, and Jon DeWitt and Andrey Gogolev for useful comments.